OpenStudy (anonymous):

A quality inspector at a valve manufacturer randomly selects one valve from each batch of fifty valves to inspect for noncompliance. The first batch of fifty valves has five non-compliant ones. The second batch of fifty valves has four non-compliant valves. Find the probability that the inspector selects a non-compliant valve both times.

OpenStudy (amistre64):

p(A) = 5/50 p(B) = 4/50 p(AnB) = p(A) * p(B) in this case